Solve x + 4y = 12 for y. Write the answer in the form y = mx + c, where m and c are fixed real numbers
Found 2 solutions by colliefan, stanbon:
Answer by colliefan(55)

Solve x + 4y = 12 for y.

x + 4y = 12
x - x +4y = 12 - x
0+4y = 12-x
4y = -x + 12
1/4 * 4y = 1/4 *(-x + 12)
1 * y = 1/4*(-x) + 1/4*(12)
y=-1/4 x +3 or y = -0.25x + 3

Answer by stanbon(48545)

Solve x + 4y = 12 for y. Write the answer in the form y = mx + c, where m and c are fixed real numbers
------------------
x+4y = 12
4y = -x + 12
y = (-1/4)x + 3
=======================
Cheers,
Stan H.